# OpticalRS-4M: Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

### v3

@@ 3.2 Efficient Pre-training
After obtaining a large-scale RS dataset, the next step is to conduct the self-supervised pre-training. However, as the capacity of the dataset grows, vast computational and time costs are required for existing MIM approaches used in the RS community, e.g., MAE [10]. To address this issue, our pipeline introduces an efficient MIM method, i.e., SelectiveMAE. In this section, we first review the preliminaries of MAE, then we present the details of SelectiveMAE.
3.2.1 Masked Autoencoders Preliminaries
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables the training of large models with reduced computational and memory requirements. 3) MAE Decoder.The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
RS optical images typically contain many redundant background pixels, which aligns well with MAE’s masking strategy. In MAE, the encoder processes only 25% of the patches, reducing computational load by avoiding processing the entire image. Building on MAE, we aim to leverage the redundancy in RS images to accelerate training. Specifically, we address two questions: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
Figure 3: Overview of SelectiveMAE. The vanilla version of SelectiveMAE.
@@ 3.2.2 Partial Reconstruction
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 3.2.1. The remaining patches serve as reconstruction targets for the decoder. Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
@@
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [45]. Experimental results in the Appendix show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
@@ 3.2.3 Progressive Semantic Token Selection
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 3.2.2. However, during training, this setup often led to issues like gradient explosions or loss divergence (see the Appendix). We speculate that, although RS images have background redundancy, using such a small portion of patches challenges MAE training. How can we achieve acceleration at a high masking ratio while ensuring MAE convergence? Curriculum learning [82, 83, 84] may offer a solution. However, it typically focuses on selecting whole image samples, whereas MIM pre-training masks large portions of image pixels, emphasizing patch-level selection. Therefore, it’s essential to distinguish between “easy” and “hard” patches in MIM pre-training.
@@
To this end, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection. In this module, depicted in Fig. 3, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
@@
For initialization, we employ a HOG selection strategy to choose the initial patch set. This simple yet effective strategy ensures that semantically rich tokens are selected. Following token initialization, we aim to incrementally increase the number of tokens to facilitate a training progression from easy to difficult, while adhering to the final masking ratio for selected tokens. This process is outlined in the Appendix. Given the initial token sets’ high HOG feature values, nearby tokens selected by PSTS also exhibit high HOG feature values. Recalling our partial reconstruction approach in Sec. 3.2.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets. Consequently, when selecting nearby tokens, the tokens input for encoding and those chosen for reconstruction are closer, thereby simplifying the training process. Conversely, if a token’s feature distribution deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn. Details of the process are presented in the Appendix.
@@ 4 Experiment
In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments. We also conducted further experiments to test the scalability of the pipeline in the ablation experiments. Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks.
Pre-training Setup. Our pre-training experiment setup largely follows MAE [10]. The masking ratio is 85%, and the reconstruction ratio is 25%. For patch selection, the adjustment threshold is set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
@@ 4.1 Ablation Study
To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-4M dataset generated by the pipeline and the accelerated training method, SelectiveMAE. Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes.
4.1.1 Superiority of OpticalRS-4M
Table 1: Comparison of downstream task performance at matched data throughput between MillionAID and OpticalRS-4M, assessed using MAE [10] pre-training and ViT Base models [39].
Dataset Method Images Number Epoch Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
MillionAID MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
OpticalRS-4M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
OpticalRS-4M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
@@ 4.1.1 Superiority of OpticalRS-4M
Dataset Diversity. In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance. We define dataset diversity as its ability to improve various types of these tasks. While datasets like MillionAID focus on classification, our OpticalRS-4M dataset is designed to extend the benefit to detection and segmentation tasks. We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput. For instance, MillionAID, with 1 million samples, is trained using the basic MAE method for 800 epochs as a benchmark. To showcase the diversity of OpticalRS-4M, containing 4 million samples, we train it with MAE for just 200 epochs. Additionally, we validate OpticalRS-4M’s diversity by progressively reducing its training data — down to 3 million, 2 million, or even 1 million samples — while keeping data throughput constant. As shown in Table 1, our experiments demonstrate that even when randomly sampling 1 million images from OpticalRS-4M, it outperforms MillionAID in downstream tasks, particularly in object detection and segmentation. This consistent superior performance, even with fewer images, highlights OpticalRS-4M’s greater diversity and effectiveness, emphasizing the benefits of using larger and more varied datasets for unsupervised pre-training.
The results in Table 1 indicate that the optimal outcomes at equivalent data throughput are achieved with 3 million samples rather than 4 million. In most cases, the model pre-trained on 3 million samples outperforms across various downstream tasks. Notably, on the DIOR and LoveDA datasets, the best results were obtained with 1 million samples trained for 800 epochs. However, as shown in Table 2, when we extended the training on 4 million samples to more epochs, the model achieved the highest average performance across most downstream tasks. This suggests that training on 4 million samples for only 200 epochs was insufficient for the MIM method to fully leverage the large dataset, resulting in inferior generalization in downstream tasks.
Table 2: Performance on downstream tasks with extended training on OpticalRS-4M.
Dataset Method Images Number Epoch Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-4M MAE 4 million 100 95.12/97.54 91.26/93.85 - - - -
OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
OpticalRS-4M MAE 4 million 400 96.36/98.16 92.41/94.03 - - - -
OpticalRS-4M MAE 4 million 800 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
Extended training on OpticalRS-4M. To maximize the potential of OpticalRS-4M, we extended training using the MAE pre-training method and ViT-B, varying only the number of training epochs while keeping other settings consistent. The fine-tuning results, shown in Table 2, reveal that as the number of epochs increases, OpticalRS-4M shows greater advantages in downstream tasks. This indicates that extended training is necessary to fully harness OpticalRS-4M’s capabilities. However, the gains diminish with additional epochs, suggesting that merely increasing epochs may not suffice. To unlock the full potential of OpticalRS-4M, exploring more effective pre-training methods, such as the proposed SelectiveMAE, may be crucial.
Table 3: Comparison of downstream task performance using SelectiveMAE and MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
Method Dataset Epoch Total Training Sence Object Semantic
Classification Detection Segmentation
Hours AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
MAE OpticalRS-4M 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
SelectiveMAE OpticalRS-4M 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
4.1.2 Effectiveness of SelectiveMAE
@@
Performance Advantage. To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-4M using both MAE [10] and SelectiveMAE. The performance of ViT-B on various downstream tasks after pre-training is presented in Table 3. The results indicate that after pre-training on OpticalRS-4M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed. This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-4M.
Table 4: The accelerating performance of SelectiveMAE for different backbones compared to MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
Model Backbone Params Total Training Hours
MAE ViT-B 86M 101
SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×)
MAE ViT-L 307M 113
SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×)
@@
Pre-training Efficiency. To further highlight the efficiency advantage of SelectiveMAE, we evaluate the pretraining time with different backbones on the MillionAID dataset, as shown in Table 4. It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger sized models.
4.1.3 Scalability of the Pipeline
@@
Table 5: Training cost (measured using 8 NVIDIA A100 GPUs) of SelectiveMAE and MAE [10] on OpticalRS-4M with different data volumes.
Method Images Number Epoch Total Training Hours
@@
To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each. The results, presented in Table 5, show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours. When the dataset was expanded to 4 million samples, the time savings increased to 81 hours. These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.

### v4

@@ 3.2 Efficient Pre-training
After obtaining a large-scale RS dataset, the next step is conduct the self-supervised pre-training. However, as the capacity of the dataset grows, vast computational and time costs are required for existing MIM approaches used in the RS community, e.g., MAE [28]. To address this issue, our pipeline introduces an efficient MIM method, SelectiveMAE. In this section, we first review the preliminaries of MAE, then we present the details of SelectiveMAE.
3.2.1 Preliminaries of MAE
Figure 3: Overview of SelectiveMAE. It inputs fewer visible patches and reconstructs only partial patches to accelerate training efficiency.
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements. 3) MAE Decoder. The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
However, applying MAE for self-supervised pre-training results in considerable training costs in terms of time complexity, especially when working with large RS datasets. To address this, we observe that RS optical images often contain redundant information, which could be omitted during training to improve pre-training efficiency. Specifically, we address two issues: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
@@ 3.2.2 Partial Reconstruction
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 3.2.1. The remaining patches serve as reconstruction targets for the decoder.
Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
@@
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [20]. Experimental results in supplementary material show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
@@ 3.2.3 Progressive Semantic Token Selection
Figure 4: Effectiveness of PSTS. Left: Using only 40% of patches for encoding and reconstruction often leads to gradient explosions. Right: The training loss after adopting PSTS.
@@
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 3.2.2. However, during training, this setup often led to issues like gradient explosions or loss divergence, as shown in Fig. 4 left. The figure shows that using an extremely low portion of patches for encoding and reconstruction caused unstable training due to gradient explosions.
@@
How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images? Inspired by the curriculum learning  [78, 7, 26], which follows the principle of learning from easy to hard, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection, as depicted in Fig. 3. In this module, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$. We define the initial set of selected token as:
[…]
We select $\lfloor\ s\times N\rfloor$ tokens with the maximum HOG feature values (i.e., ${top}_{\lfloor\ s\times N\rfloor}$) from the original token set to form the initial token set. This simple yet effective strategy ensures that semantically rich tokens are selected.
@@
After token initialization, we incrementally increase the number of tokens to guide training from easier to more challenging examples, while maintaining the final masking ratio for selected tokens, as outlined in Algorithm 1. Given the high HOG feature values of the initial token sets, nearby tokens selected by PSTS also exhibit high HOG values. In partial reconstruction method in Section 3.2.2, we filter unselected tokens to retain those with high HOG values as reconstruction targets. Thus, selecting nearby tokens brings encoding tokens and reconstruction targets closer, streamlining training. In contrast, if a token’s feature distribution differs significantly from others, we consider these more challenging patches to learn.
Algorithm 1 Progressive Semantic Token Selection
1:Number of training epochs $T$ and total training stages $N_{g}$ (i.e., $T/N_{g}$ epochs for each stage), masking rate $m$, input dataset $\mathcal{X}$
2:Obtain the selected tokens set ${S^{K}}$ and update ${S^{U}}$ in each epoch
3:for $t\leftarrow 1$ to $T$ do
4:    Sample data sample from $\mathcal{X}$, feed-forwarded through the embedding to obtain the output token set ${S}_{N}$
5:    $s\leftarrow\frac{1}{2}(1-m)$
6:    Obtain ${S}^{I}$ via Eq. (2) and initialize ${S^{U}}$
7:    Obtain the current training stage $\zeta=\lceil N_{g}*t/T\rceil$
[…]
9:    Obtain $S^{*}$ and update $S^{K},S^{U}$ via Eq. (5) and Eq. (6)
10:end for
Specifically, we select tokens from $S^{U}$ based on $S^{I}$. First, we use $\mathbb{S^{I}}\in\mathbb{R}^{|S^{K}|\times d}$ and $\mathbb{S^{U}}\in\mathbb{R}^{|S^{U}|\times d}$ to denote the matrix representation of the initial token set $S^{I}$ and the unselected token set $S^{U}$, where ||$\cdot$|| ⋅ | represents the number of tokens and $d$ the feature dimension after the embedding layer. We use Cosine Distance to measure the dissimilarity between the tokens in these two sets:
$\small\mathcal{D}(S^{U},S^{I})=\mathbf{1}-\cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-0.96873pt\left|\mathbb{S^{U}}\right|\kern-0.96873pt\right|\cdot\left|\kern-0.96873pt\left|\mathbb{S^{I}}\right|\kern-0.96873pt\right|},$ (3)
where $\mathbf{1}$ is an all-one matrix. $\mathcal{D}(S^{U},S^{I})\in\mathbb{R}^{|S^{U}|\times|S^{I}|}$ represents the pairwise distances between tokens in $S^{U}$ and $S^{I}$. Next, we define the distance between the tokens in $S^{U}$ to the initial token set $S^{K}$ based on the selection criteria in each training stage as follows:
[…]
where $i\in\{1,\cdots,|S^{U}|\}$, $j\in\{1,\cdots,|S^{I}|\}$, and $\zeta$ represents the training stage depending on the number of epochs. Finally, we sample $\lfloor N\times(1-m-s)\rfloor$ tokens from ${S^{U}}$ and add them with ${S^{I}}$ to form ${S^{K}}$, which can be formulated as follows:
[…]
$\small{S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$. The operations in Eq. (5) and Eq. (6) are performed in each training epoch. This process is summarized in Algorithm 1. Fig. 5 further provides an example to illustrate which patches should be selected at different stages, helping to facilitate readers’ understanding.
Figure 5: Illustration of the patch selection process in PSTS.
@@ 4 Experiment
Table 1: Performance comparison of using the MillionAID and OpticalRS-13M datasets separately with the same data throughput, evaluated by MAE [28] and ViT-B [19].
Dataset Method Image Number Epoch Scene Object Semantic
Classification Detection Segmentation
AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
MillionAID [48] MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
OpticalRS-13M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
OpticalRS-13M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
OpticalRS-13M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-13M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
OpticalRS-13M MAE 8 million 100 96.58/98.26 91.83/93.99 75.20 66.99 52.56 79.43
OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60 74.50 66.36 52.63 79.39
In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments. We also conducted further experiments to test the scalability of the pipeline in the ablation experiments. Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks. All experiments are performed by fine-tuning pre-trained models on downstream tasks, including RS scene classification, object detection, and semantic segmentation. The pre-training and fine-tuning settings of our methods can be found in the supplementary material.
Table 2: The performance comparison between SelectiveMAE and the baseline method MAE [28] with the ViT-B backbone. The training time is measured on NVIDIA A100 GPUs. To accelerate validation, we select 4 million images from OpticalRS-13M for pre-training.
Method Image Number Epoch Total Training Scene Object Semantic
Classification Detection Segmentation
Time (h) AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
MAE 4 million 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
SelectiveMAE 4 million 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
@@ 4.1 Ablation Study
To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-13M dataset generated by the pipeline and the accelerated training method, SelectiveMAE. Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes. More ablation experiment results related to model design are presented in the supplementary material.
@@
Dataset Diversity of OpticalRS-13M. In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance. Here, We define dataset diversity as the ability to improve performance across various types of tasks. While datasets like MillionAID [48] focus on classification, our OpticalRS-13M dataset is designed to extend the benefit to detection and segmentation tasks. We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput. As shown in Table 1, even with random sampling of 1 million images from OpticalRS-13M, it outperforms MillionAID on downstream tasks, particularly in object detection and segmentation. The results in Table 1 also indicate that the optimal outcomes at equivalent data throughput are achieved by different settings other than using all 13 million images. This suggests that training on a large number of samples with few epochs was insufficient for the MIM method to fully leverage the large dataset, leading to inferior generalization on downstream tasks. Experimental results in the supplementary material show consistent performance improvements across downstream tasks with more training epochs.
Table 3: The efficiency comparison between SelectiveMAE and the baseline method MAE [28] across different backbones on the MillionAID dataset [48] with 1600 epochs. The memory is measured on a single NVIDIA A100 GPU with a batch size of 256.
Model Backbone Params Total Training Time (h) GPU Memory (MB)
MAE ViT-B 86M 101 17628
SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×) 9570 (1.8×1.8$\times$1.8 ×)
MAE ViT-L 307M 113 30530
SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×) 18790 (1.6×1.6$\times$1.6 ×)
@@
Performance Advantage of SelectiveMAE. To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-13M using both MAE [28] and SelectiveMAE. The performance of ViT-B on various downstream tasks after pre-training is presented in Table 2. The results indicate that after pre-training on OpticalRS-13M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed. This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-13M.
@@
Efficiency Advantage of SelectiveMAE. To further highlight the efficiency advantage of SelectiveMAE, we evaluate the training time and memory footprint during pretraining of different backbones on the MillionAID dataset, as shown in Table 3 (the corresponding accuracies are presented in Table 5). It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled training acceleration compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger models. Additionally, it is worth noting that SelectiveMAE also reduces memory footprint by using fewer patches.
Scalability of the Pipeline To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each. The results in Table 4 show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours. When the dataset was expanded to 4 million samples, the time savings increased to 81 hours. These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
@@
Table 4: The training cost of SelectiveMAE and the baseline method MAE [28] on OpticalRS-13M with different data volumes. All experiments are conducted on NVIDIA A100 GPUs.
Method Images Number Epoch Total Training Time (h)
@@
Table 5: Performance comparison results of different models. “TR” represents the ratio of training data to the entire dataset. The first and second scores are marked by bold and blue, respectively. * means pre-training on 4 million images sampled from OpticalRS-13M.
Model Backbone Params (M) Scene Object Semantic
Classification Detection Segmentation
AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
SeCo [49] ResNet-50 26 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [1] ResNet-50 26 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [65] ResNet-50 26 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [50] ResNet-50 26 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [17] ViT-L 307 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [56] ViT-L 307 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [79] ViT-S 22 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [63] Swin-B 88 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [5] Swin-B 88 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [51] Swin-B 88 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [70] ViT-B+RVSA 86 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
LSKNet [44] LSKNet-S 14 97.05/98.22 94.27/95.83 - - 54.00 -
OREOLE [18] ViT-G 914 96.71/   -    -   /   - 77.40 71.31 54.00 -
SelectiveMAE* ViT-B 86 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
SelectiveMAE* ViT-L 307 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
SelectiveMAE ViT-B 86 96.64/98.30 92.53/94.50 75.10 67.69 52.68 79.44
SelectiveMAE ViT-L 307 97.21/98.48 94.16/95.90 78.70 71.75 53.92 79.48